A copper road whose length at 70°c is 4.0m is heated to 10°c find it

1
Expert's answer
2022-02-20T08:20:02-0500

Given:

"l_0=4.0\;\rm m"

"\Delta T=30\:\rm K"

"\alpha=17*10^{-6}\:\rm K^{-1}"


The thermal linear expansion is described by an equation

"\Delta l=l_0\alpha\Delta T\\
=\rm4.0\: m*17*10^{-6}\:\rm K^{-1}*30\: K=0.002\: m=2\: mm"